What is the value of k with perpendicular equations 24x 4y=-5 and kx-3y=0

Answer by nanakwame(5)

24x 4y=-5~1
kx-3y=0~2
gradient of eqn1=6
gradient of eqn2=k/3
6=-3/k
:k=-1/2